### v1

# InsFusion: Rethink Instance-level LiDAR-Camera Fusion for 3D Object Detection

Abstract
Three-dimensional Object Detection from multi-view cameras and LiDAR is a crucial component for autonomous driving and smart transportation. However, in the process of basic feature extraction, perspective transformation, and feature fusion, noise and error will gradually accumulate. To address this issue, we propose InsFusion, which can extract proposals from both raw and fused features and utilizes these proposals to query the raw features, thereby mitigating the impact of accumulated errors. Additionally, by incorporating attention mechanisms applied to the raw features, it thereby mitigates the impact of accumulated errors. Experiments on the nuScenes dataset demonstrate that InsFusion is compatible with various advanced baseline methods and delivers new state-of-the-art performance for 3D object detection.
1 Introduction
Figure 1: Comparison between the existing paradigm and our paradigm.
Multi-sensor fusion is essential for an accurate and reliable perception system, where the workflow typically entails extracting raw features from images and point clouds, transforming the 2D features into 3D representations, converting the 3D features from each modality into a consistent coordinate frame, fusing them, and extracting instance features from multimodal features. At each step of this process, various forms of noise or errors can be introduced. For instance, these may include depth estimation errors during 2D-to-3D transformation or offset prediction inaccuracies in deformable attention mechanisms, inaccuracies in extrinsic parameters during coordinate transformation, as well as information loss during the fusion process. These errors can accumulate and spread, ultimately affecting overall performance.
To address this issue, we introduce InsFusion, a universal instance-level fusion paradigm specifically tailored for Bird’s-Eye View (BEV) LiDAR-camera fusion models. InsFusion extracts proposals from both raw and fused features and utilizes these proposals to query the raw features, thereby alleviating the impact of accumulated errors. InsFusion can be easily integrated into all existing BEV-based LiDAR-camera fusion models, enhancing their performance.
Our contributions are: (1) We propose InsFusion, a novel paradigm compatible with existing multimodal 3D object detection methods, to address the issue of accumulated errors. (2) InsFusion requires only minimal fine-tuning of the baseline models, resulting in low training cost. (3) InsFusion enhances the performance of both advanced models chen2023focalformer3d ; yin2024fusion and achieves state-of-the-art results on the nuScenes dataset Caesar2020nuScenesAM .
2 Related Work
LiDAR-only. Prior LiDAR-based works directly operate on the raw LiDAR point clouds qi2017pointnet++ ; Qi2017PointNetDL ; qi2018frustumPF ; Shi2019PointRCNN3O ; Yang20203DSSDP3 ; li2021lidar to extract features. With the growing volume of LiDAR points in outdoor datasets, several works transform point clouds into Euclidean feature space, such as 3D voxels Zhou2018VoxelNetEL , range images fan2021rangedet ; Sun2021RSNRS , and bird’s eye view (BEV) plane Lang2019PointPillarsFE ; Wang2020PillarbasedOD ; Yin2020Centerbased3O ; fan2021rangedet ; Sun2021RSNRS .
Camera-only. Early works lu2021geometry ; Roddick2019OrthographicFT ; liu2021autoshape ; kumar2021groomed ; zhang2021objects ; zhou2021monocular ; reading2021categorical ; wang2021progressive ; wang2021depth predicted 3D bounding boxes from monocular images. With the introduction of datasets such as nuScenes Caesar2020nuScenesAM , methods for multi-view cameras have demonstrated superior performance. The core problem they address involves transforming 2D feature maps into 3D features while extracting instance features. LSS philion2020lift and related methods huang2021bevdet ; li2023bevdepth ; sts ; HoP ; SOLOFusion ; xia2024henet employ depth estimation to lift 2D features into the BEV space, subsequently extracting instance features through techniques such as heatmap peak detection Yin2020Centerbased3O . BEVFormer series bevformer ; bevformerv2 acquires BEV features through Deformable Attention zhu2020deformable . Subsequent approaches wang2022detr3d ; liu2022petr ; liu2023petrv2 ; sparse4d ; sparse4dv2 ; wang2023exploring ; liu2023sparsebev directly obtain instance features by querying 2D feature maps. They streamline the feature extraction pipeline, thereby mitigating cumulative errors and consequently achieving superior performance in object detection tasks.
LiDAR-camera fusion. In contrast to early-fusion Huang2020EPNetEP ; Sindagi2019MVXNet ; Wang2021PointAugmentingCA ; Yin2021MVP or late-fusion pang2020clocs strategies, feature-level fusion methods have demonstrated superior performance. Existing feature-level fusion methods Yoo20203DCVFGJ ; bai2022transfusion ; liang2022bevfusion ; liu2022bevfusion are primarily extensions of LiDAR-only approaches, with BEV-based methods constituting the predominant paradigm. Such paradigms typically cascade multiple modules, thereby introducing cumulative errors. Subsequent methods often aim to mitigate inaccuracies in specific components of this pipeline, such as reducing depth estimation errors hu2023ea ; zhao2024simplebev to alleviate noise in the transformation from 2D feature maps to BEV features. IS-Fusion yin2024fusion employs instance-guided Fusion to mitigate noise introduction during the BEV fusion process. FocalFormer3D chen2023focalformer3d uses multistage heatmaps to reduce errors in the instance feature extraction process from multimodal BEV features. In contrast to previous efforts that primarily focus on refining individual components within the pipeline, InsFusion leverages minimally processed raw features to mitigate noise while maintaining full compatibility with existing methods, thereby enhancing their performance.
3 Methodology
As shown in Fig. 2, in order to extract information from features with minor cumulative noise, InsFusion takes three steps: extracting proposals from 2D feature maps, LiDAR features, and fused features; performing query alignment; and refining instance features through attention mechanisms.
Figure 2: Overview of the InsFusion framework. The framework extracts proposals from raw camera features, LiDAR features, as well as fused BEV features, and then aligns and refines all proposals to predict 3D bounding boxes.
3.1 Extract Queries From Raw Features
Camera Branch. For image features $F_{img}\in\mathbb{R}^{h\times w\times C_{img}}$, we first define $K$ learnable camera queries, which are drawn from random Gaussian distributions and attached to a $D_{q}$-dim query feature to encode the rich instance characteristics. Then we adopt adaptive sampling and mixing liu2023sparsebev to obtain instance features $Q_{img}^{(0)}\in\mathbb{R}^{K\times D_{q}}$.
Lidar Branch. For LiDAR BEV features $F_{lidar\_bev}\in R^{X\times Y\times C_{lidar\_bev}}$, InsFusion predicts the instance heatmap and extracts proposals through peak detection. We select the highest K peaks from $H_{\text{lidar}}$ and encode their spatial coordinates into initial embeddings, forming the initial LiDAR query $Q_{\text{lidar}}^{(0)}\in\mathbb{R}^{K\times D_{q}}$.
Fusion branch. InsFusion is a framework compatible with existing multimodal fusion methods, applicable to any feature fusion strategy, whether it employs BEV heatmaps or Transformer-based architectures to obtain the fused instance features $Q_{\text{fusion}}^{(0)}$.
Queries Alignment. In this stage, the three sets of queries ($Q_{\text{img}}\in\mathbb{R}^{K\times D_{q}}$, $Q_{\text{lidar}}\in\mathbb{R}^{K\times D_{q}}$, $Q_{\text{fusion}}\in\mathbb{R}^{K\times D_{q}}$) are projected into a shared latent space via modality-specific linear transformers, eliminating distribution shifts while preserving modality-specific strengths: $\hat{Q}_{\text{img}}=W_{\text{img}}\cdot Q_{\text{img}}+b_{\text{img}}$, $\hat{Q}_{\text{lidar}}=W_{\text{lidar}}\cdot Q_{\text{lidar}}+b_{\text{lidar}}$, $\hat{Q}_{\text{fusion}}=W_{\text{fusion}}\cdot Q_{\text{fusion}}+b_{\text{fusion}}$, where $W_{\text{img}},W_{\text{lidar}},W_{\text{fusion}}\in\mathbb{R}^{D_{q}\times D_{q}}$ are learnable projection matrices, and $b_{\text{img}},b_{\text{lidar}},b_{\text{fusion}}\in\mathbb{R}^{1\times D_{q}}$ are bias terms. These parameters are optimized end-to-end with the entire network.
3.2 Instance Feature Refinement
After alignment, we concatenate the three aligned queries sets ($\hat{Q}_{\text{img}},\hat{Q}_{\text{lidar}},\hat{Q}_{\text{fusion}}$) to form a comprehensive query $Q^{(0)}$. This design integrates low-error information from raw-modality proposals and cross-modal context from fusion proposals, enabling multi-source feature interaction during refinement. Instead of using a single feature map as Key (K) and Value (V), we leverage the three core feature sources (raw camera features $F_{\text{img}}$, raw LiDAR BEV features $F_{lidar\_bev}$, and fused BEV features $F_{fusion_{b}ev}$) as separate Key-Value. The concatenated query $Q_{\text{concat}}$ performs targeted queries on each of these three feature sources via a deformable transformer decoder—this ensures the query pool fully captures modality-specific strengths (e.g., image semantics, LiDAR geometry, cross-modal complementarity) while avoiding error accumulation from over-reliance on a single feature source. The refinement process follows:
$Q^{(0)}=\left[\hat{Q}{\text{img}},\hat{Q}{\text{lidar}},\hat{Q}{\text{fusion}}\right]\in\mathbb{R}^{3K\times D_{q}},$ (1)
$Q^{(l)}=\text{DeformableTransformerLayer}(Q^{(l-1)},\text{Flatten}(F_{\text{img, raw}}),\text{Flatten}(F_{\text{lidar, raw}}),\text{Flatten}(F_{\text{fusion}})).$
This process is repeated for $L$ iterations. Here, $\left[\cdot,\cdot,\cdot\right]$ denotes concatenation, and the deformable decoder layer independently computes attention between $Q_{\text{concat}}^{(l-1)}$ and each of the three flattened Key-Value pairs (i.e., $F_{img}$, $F_{lidar_{b}ev}$, $F_{fusion\_bev}$). The attention outputs from the three feature sources are aggregated via element-wise addition to update the query. This multi-source query design ensures InsFusion fully leverages the precision of raw features and the complementarity of fused features, directly addressing error accumulation by anchoring refinement to low-error, modality-specific sources.
4 Experiments
Table 1: Results on the nuScenes val set. Applying InsFusion to FocalFormer and the state-of-the-art method IS-Fusion effectively enhances their performance.
Method Mod. I.B. mAP NDS Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
BEVFormer bevformer C V2-99 41.6 51.7 61.8 37.0 12.8 44.4 17.2 52.5 42.9 39.8 49.4 58.4
BEVDet4D huang2021bevdet C R50 40.8 52.3 62.7 33.4 12.4 39.3 15.6 58.7 38.5 39.1 46.7 61.8
SparseBEV liu2023sparsebev C V2-99 57.3 65.0 72.5 53.8 25.9 61.3 33.3 67.7 58.9 60.8 65.2 73.9
CenterPoint Yin2020Centerbased3O L - 58.6 65.8 84.0 53.8 19.9 68.1 38.7 67.8 62.2 43.4 83.1 65.3
Transfusion-L bai2022transfusion L - 65.1 70.1 86.6 41.1 26.1 70.6 38.1 68.5 69.4 54.7 85.5 71.9
FocalFormer3D-L chen2023focalformer3d L - 66.3 70.1 87.4 59.1 27.8 76.4 44.3 71.1 75.3 60.4 87.2 74.0
FUTR3D chen2022futr3d L+C R101 64.2 68.0 86.3 61.5 26.0 71.9 42.1 64.4 73.6 63.3 82.6 70.1
BEVFusion liang2022bevfusion L+C Swin-T 69.6 72.1 89.1 66.7 30.9 77.7 42.6 73.5 79.0 67.5 89.4 79.3
FocalFormer3D chen2023focalformer3d L+C R50 70.5 73.1 89.4 65.4 30.5 78.8 45.6 73.2 83.0 70.9 88.7 79.5
+InsFusion (ours) L+C R50 71.5(+1.0) 74.2(+1.1) 88.9 69.7 30.5 77.7 53.1 84.2 78.4 65.0 89.0 78.9
IS-Fusion yin2024fusion L+C Swin-T 72.3 73.7 89.4 68.3 36.1 79.4 51.6 74.7 80.6 70.1 89.3 83.6
+InsFusion (ours) L+C Swin-T 73.4(+1.1) 74.3(+0.6) 89.4 68.1 37.0 79.5 52.9 80.1 80.3 70.4 91.3 85.4
’Mod.’: Modality, ’I.B.’: Image Backbone. ‡ : Reproduced with official code. Notion of modality: Camera (C), LiDAR (L).
Notion of class: Construction vehicle (C.V.), pedestrian (Ped.), traffic cone (T.C.).
We provide experimental setup and implementation details in the Appendix  A.
Main Result. Experimental results on the nuScenes Caesar2020nuScenesAM dataset are summarized in Tab. 1. When integrated with FocalFormer3D, InsFusion yields performance gains of +1.0% mAP and +1.1% NDS; when applied to the state-of-the-art method IS-Fusion, InsFusion achieves improvements of +1.1% mAP and +0.6% NDS. These results show the effectiveness of InsFusion. Unfortunately, due to a service outage in the official nuScenes evaluation server, we were unable to evaluate the results on the test set.
Ablation study for the number of Transformer layers. As shown in Tab. 2, when $L=1$, the single-layer deformable Transformer exhibits insufficient capability to refine concatenated proposals, which prevents the full capture of complementary information across multi-source features (raw camera, raw LiDAR, and fused BEV); and when $L=6$, excessive iterative refinement induces the model to overfit to local noise generated during the fusion process rather than focusing on critical global instance-level information, thereby leading to performance degradation, $L=2$ is ultimately identified as the optimal configuration, as it achieves a balanced trade-off between ensuring adequate feature refinement and avoiding overfitting to local noise.
Table 2: Ablation study on the number of deformable transformer layers. The metrics mAP (%) and NDS (%) are computed on the nuScenes val set.
Layers Number ($L$) mAP (%) NDS (%)
1 67.31 71.35
2 69.33 72.35
6 68.82 71.59
5 Conclusion
3D object detection using multi-view cameras and LiDAR is crucial for autonomous driving and smart transportation. However, it suffers from noise and error accumulation during feature processing and fusion, which degrades performance. To solve this, we proposed InsFusion, a universal instance-level fusion paradigm for LiDAR-camera fusion. It extracts queries from both raw and fused features, and then queries these features to alleviate accumulated errors. InsFusion is compatible with existing BEV-based fusion models and only requires minimal fine-tuning (with low training cost). Experiments on the nuScenes dataset demonstrate that InsFusion can further enhance the performance of several high-performing baseline methods.
Appendix A Experimental Setup.
Dataset. Our experimental evaluation is performed on the nuScenes dataset, a standard benchmark for autonomous driving research that provides comprehensive multi-sensor data across diverse urban environments. This dataset enables rigorous evaluation of 3D detection systems under challenging real-world conditions.We adopt the official evaluation metrics: NuScenes Detection Score (NDS) and mean Average Precision (mAP), which comprehensively assess detection accuracy, localization precision, and orientation prediction.
Implementation Details. Our main implementation relies on the open-source MMdetection3D [54] framework. When deploying InsFusion on the FocalFormer and IS-Fusion baselines, we maintain all the original settings of the baseline model, including the LiDAR backbone, camera backbone, and the original fusion network, to ensure a fair comparison.
For the key parameters of InsFusion: the number of layers for the Shared Deformable Transformer is set to $L=2$; the initial query count for both the LiDAR branch and camera branch is 300. We train our network using a two-stage training mode: Firstly, train the baseline model (skip this step if official weights are available and use the official weights directly), and then train the "extract proposals from img" network. When training the InsFusion-enhanced model, load and freeze the baseline weights, and load the weights of the "extract proposals from img" network.
For optimization, we use the Adam optimizer with a one-cycle learning rate policy (max learning rate $2\times 10^{-5}$) and a weight decay of 0.01. All experiments are conducted on 8 NVIDIA RTX 8000 GPUs, with a batch size of 16 and a training duration of 6 epochs.
Appendix B Limitation.
A limitation of InsFusion is that it introduces additional computational overhead compared to baseline models, which manifests as a slight reduction in inference efficiency. However, as shown in Tab. 3, this efficiency cost is insignificant and well within acceptable limits for practical use: when integrated into the FocalFormer baseline, the inference FPS decreases by only 9.3%; when applied to the IS-Fusion baseline, the FPS reduction is even smaller, at just 6.6%. Based on the inference FPS metric alone, the resulting decrease in inference latency remains within an acceptable range, indicating that its computational overhead does not impede practical deployment in 3D detection applications.
Table 3: Computational Overhead Comparison Between InsFusion-Enhanced Models and Baselines
[…]
FocalFormer 1.29
+InsFusion 1.17
IS-Fusion 0.91
+InsFusion 0.85
Appendix C Broader Impacts Statement.
All datasets we used are published datasets. We do not see potential privacy-related issues. This study may inspire future research on LiDAR-camera fusion model for 3D perception.